OpenStudy (anonymous):

Multiply: \[(3\sqrt{3}-6\sqrt{2})(5\sqrt{3}+10\sqrt{2})\]

OpenStudy (saifoo.khan):

\[\Huge -75\]

jhonyy9 (jhonyy9):

3sqrt3*5sqrt3 + 3sqrt3*10sqrt2 - 30sqrt6 - 6sqrt2*10sqrt2 15*3 +30sqrt6 -30sqrt6 -60*2 45-120 -75

OpenStudy (anonymous):

Multiply and Simplify by factoring. Assume that all expressions under radicals represent nonnegative numbers: \[\sqrt[3]{y ^{10}} \sqrt[3]{16y ^{11}}\]

OpenStudy (saifoo.khan):

i have dont that before!! :D

OpenStudy (saifoo.khan):

Done*

OpenStudy (saifoo.khan):

\[\Huge 2y^7\sqrt[3]{2}\]

OpenStudy (anonymous):

Multiply and Simplify by factoring. Assume that all expressions under radicals represent nonnegative numbers

OpenStudy (anonymous):

\[\sqrt[3]{y ^{4}} \sqrt[3]{81y ^{5}}\]